OpenStudy (anonymous):

Find the cube roots of 8(cos 216° + i sin 216°).

OpenStudy (anonymous):

@jagatuba

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

What math is this?

OpenStudy (anonymous):

Pre-Cal

OpenStudy (luigi0210):

I don't think I've ever done something like this..

OpenStudy (loser66):

r = 8 (cos216 + isin 216) \(r^{1/3} = 8^{1/3}(cos (\dfrac{216}{3})+isin(\dfrac{216}{3})\) = \(2(cos(72) + isin (72))\)

OpenStudy (loser66):

That's called De Moivre's theorem

OpenStudy (anonymous):

okay thanks for explaining it to me!I see you plug it in for me as well thanks for that!